Kathy Stonewall bought a new car for ​$14 comma 714. A​ dealer's financing was available through a local bank at an interest rate of 13.5​% compounded monthly. Dealer financing required a 8​% down payment and 48 equal monthly payments. Because the interest rate was rather​ high, Kathy checked her credit union for possible financing. The loan officer at the credit union quoted a 11.8​% interest rate for a​ new-car loan and 12.5​% for a used car. But to be eligible for the​ loan, Kathy has to be a member of the union for at least six months. Since she joined the union two months​ ago, she has to wait four more months to apply for the loan.​ Consequently, she decided to go ahead with the​ dealer's financing, and four months later she refinanced the balance through the credit union at an interest rate of 12.5​%. ​(a) Compute the monthly payment to the dealer. The monthly payment to the dealer was ​$ nothing.​ (Round to the nearest​ cent.) ​(b) Compute the monthly payment to the union. The monthly payment to the union was ​$ nothing.​ (Round to the nearest​ cent.) ​(c) What is the total interest payment on each​ loan? The total interest payment to the dealer was ​$ nothing.​ (Round to the nearest​ cent.)
Answer #1

Using Excel

Value of new car 14714
Downpayment 8%
Downpayment amount 1177.12
Loan required 13536.88
Dealer interest rate 13.50%
Dealer interest rate (per month) 1.125%
Dealer loan ternure (months) 48
a Monthlty payments 366.53
b Loan installments left after 4 payment 44
Loan due 12,665.35
interrst rate charged by union 12.50%
Monthly payment to union 360.31
c Total interest paid to dealer 594.59
Total amount paid to union 15,853.49
Total interest paid to union 3,188.14

Showing formula in excel

Value of new car 14714
Downpayment 0.08
Downpayment amount =F8*F7
Loan required =F7-F9
Dealer interest rate 0.135
Dealer interest rate (per month) =F11/12
Dealer loan ternure (months) 48
a Monthlty payments =PMT(F12,F13,-F10)
b Loan installments left after 4 payment =F13-4
Loan due =PV(F12,F15,-F14)
interrst rate charged by union 0.125
Monthly payment to union =PMT(F17/12,F15,-F16)
c Total interest paid to dealer =IPMT(F12,1,F13,-F10) + IPMT(F12,2,F13,-F10)+IPMT(F12,3,F13,-F10)+IPMT(F12,4,F13,-F10)
Total amount paid to union =F18*F15
Total interest paid to union =F20-F16